OpenStudy (anonymous):

What is the expression for the sum\[1+3x+5x^{2}+7x^{3}+9x^{4}....\]

OpenStudy (anonymous):

there is an expression,but only if we are given x is a fraction..is it in this case ?

hartnn (hartnn):

is x<1 ?

OpenStudy (anonymous):

there is no criteria

OpenStudy (anonymous):

then this cant be simplified to some expression i'd guess.. you see,,the derivation of the expression assumed x<1..

OpenStudy (anonymous):

ok let x < 1

OpenStudy (anonymous):

then the expression of sum is ab/(1-r) + (dbr)/( (1-r)^2 ) a = 1st term of AP b= 1st term of GP d & r have there usual meanings..

hartnn (hartnn):

this is arithmetico geometric series..................

OpenStudy (anonymous):

the above thing has a very simple yet interesting derivation/proof..

OpenStudy (anonymous):

1+3x+5x^2+7x^3 +9x^4... =1+x+2x+x^2+4x^2 +x^3+6x^3+x^4+8x^4+... =(1+x+x^2+x^3+x^4+...)+(2x+4x^2+6x^3+8x^4+...) =(x^0+x^1+x^2+x^4+...)+2(x+2x^2+3x^3+4x^4+...)

OpenStudy (anonymous):

That may lead to the solutiom

OpenStudy (anonymous):

I got this answer\[\frac{ 3-x-2x^{n} }{ (x-1)^{2} }\] is it right?

OpenStudy (anonymous):

Thanks